OpenStudy (anonymous):

A narrow beam of monochromatic light is incident normally on diffraction grating. 3rd order diffraaacted beams are formed at angle 45 to the original direction. Whats the highest order of difracted beam produced by this grating

OpenStudy (anonymous):

You need to use the equation \[d \sin \theta = n \lambda \]. The angle \[\theta \] is 45 degrees when n = 3. You need more information to answer this question ... like the wavelength and the spacing between slits in the grating ..